5. A boat is sailing along a circular path of v= (0.0625t^2) m⁄s where t is in seconds. Determine the magnitude of its acceleration when t= 10 s.
1
Expert's answer
2020-08-21T11:02:32-0400

"v = 0.0625t^2"

The acceleration by definition is "a(t) = \frac{dv}{dt}= 0.0625 \cdot 2 \cdot t= 0.125t"

At the time t=10 s: "a(10) = 0.125 \cdot 10= 1.25 \; \frac{m}{s^2}"

Answer: "a(10) = 1.25 \; \frac{m}{s^2}"